OpenStudy (anonymous):

simplify: n^2+3n+2/2 divided by n+1/2(n+2)

OpenStudy (callisto):

\[ \large \frac{\frac{n^2+3n+2}{2}}{\frac{n+1}{2(n+2)}}\]\[=\large \frac{\frac{n^2+n+2n+2}{2}}{\frac{n+1}{2(n+2)}}\]\[=\large \frac{\frac{n(n+1)+2(n+1)}{2}}{\frac{n+1}{2(n+2)}}\]\[=\large \frac{\frac{(n+2)(n+1)}{2}}{\frac{n+1}{2(n+2)}}\]\[=\large \frac{(n+2)(n+1)}{2}\times\frac{2(n+2)}{n+1}\] Now cancel the common factors and simplify it.